6. A small indoor green/:iouse (herbarium ) is made of glass panes (including base) held t.ogether with tape. It is 30 cm wng, 26 cm wU:ie and 26 cm high.
(i) What is the area of the glass?
(ii) How much of tape is needed for all the 12 edges?

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
i) Area of glass (total surface area) = 2(30 x 25 + 25 x 25 + 30 x 25) cm.2 = 2(750 + 625 + 750) cm2 = 4250 cm2. (ii) 1btal tap needed = 4(l + b + h) = 4(30 + 25 + 25) cm = 320 cm.